three years ago adams age was nine times that of judy.what is the combined age of adam and judy when judy's age is half of adam?
answer choices: a) 17
b) 26
c) 24
d) 45

Answer by ankor@dixie-net.com(22740)


It says "Three years ago adams age was nine times that of judy."
:
Let's say 3 years ago: Judy's age = J; and Adams's age = 9J
:
what is the combined age of adam and judy when judy's age is half of adam?
Another words in x number of years: A's age is twice J's age:
:
Adam is twice Judy's age
9J + x = 2(J + x)
9J + x = 2J + 2x
9J - 2J = 2x - x
7J = x
Try this if J's age = 1, Then x = 7 years. We know A's age is 9
In 7 years J = 8, A = 16, There combined ages would be 24
:
Check:
If J = 2 then x = 14, A = 18, in 14 years then J = 16 and A = 32, their
combined age of 48 is not one of the answers
:
Not a very neat solution, hope it helps